A vendor buys plastic items for Rs. 540 a d sells them at a loss of 9 percent. Find the selling price.
Answer by Theo(13342)

his loss is 9% and his cost is 540.

therefore 540 - .09 * 540 = the selling price.

solve for the selling price to get selling price = 540 - 48.6 = 491.4

his cost is 540.
his selling price is 491.4
his loss is 540 - 491.4 = 48.6
48.6 / 540 = .09 = 9%.